### v1

@@ Abstract
Reinforcement learning agents can achieve superhuman performance, but their decisions are often difficult to interpret. This lack of transparency limits deployment, especially in safety-critical settings where human trust and accountability are essential. In this work, we develop a theoretical framework for explaining reinforcement learning through the influence of state features—what the agent observes in its environment. We identify three core elements of the agent-environment interaction that benefit from explanation: behaviour (what the agent does), performance (what the agent achieves), and value estimation (what the agent expects to achieve). We treat state features as players cooperating to produce each element and apply Shapley values—a principled method from cooperative game theory—to identify the influence of each feature. This approach yields a family of mathematically grounded explanations with clear semantics and theoretical guarantees. We use illustrative examples to show how these explanations align with human intuition and reveal novel insights. Our framework unifies and extends prior work, making explicit the assumptions behind existing approaches, and offers a principled foundation for more interpretable and trustworthy reinforcement learning.
Keywords: Reinforcement learning, Shapley values, explainable artificial intelligence, explainable reinforcement learning, Markov decision process, feature-based explanations
@@ 1 Introduction
Reinforcement learning is an approach to artificial intelligence that enables agents to learn through trial-and-error interaction with their environment. By maximising a scalar reward signal that provides feedback on their actions, agents learn to achieve desired outcomes. Reinforcement learning has achieved remarkable success across diverse domains, from mastering complex games (Mnih et al., 2015; Silver et al., 2016, 2017; Vinyals et al., 2019; Schrittwieser et al., 2020; Wurman et al., 2022) to controlling real-world systems (Bellemare et al., 2020; Degrave et al., 2022; Seo et al., 2024). Unlike humans, reinforcement learning agents typically cannot explain their behaviour, making it difficult for people to understand or trust their decisions. This hinders deployment, especially in safety-critical domains where human accountability is essential. For example, if a human operator could not understand a reinforcement learning agent’s decisions, would they deploy it to regulate a nuclear fusion reactor (Degrave et al., 2022; Seo et al., 2024)?
In this work, we address this limitation by introducing a theoretical framework for explaining reinforcement learning. We identify three core elements of reinforcement learning that benefit from explanation: behaviour (what the agent does), performance (what reward the agent achieves), and value estimation (what reward the agent expects to achieve). Together, these elements offer a comprehensive understanding of how agents interact with their environments.
We propose to explain these three elements through the influence of state features—numerical values that describe what the agent observes in its environment. By treating features as players cooperating to produce behaviour, performance, or value estimates, we cast the problem in terms of cooperative game theory. This framing allows us to attribute each feature’s contribution using Shapley values (Shapley, 1953)—a principled method for fairly distributing credit among contributors—with the resulting explanations uniquely characterised by a set of intuitive mathematical axioms. The outcome is a theoretically principled framework: Shapley Values for Explaining Reinforcement Learning (SVERL). Using illustrative examples, we show that SVERL produces explanations that align with human intuition while revealing novel insights into an agent’s behaviour, outcomes, and expectations.
Although Shapley values have been used in reinforcement learning (Rizzo et al., 2019; Carbone, 2020; He et al., 2021; Wang et al., 2020; Zhang et al., 2020; Liessner et al., 2021; Løver et al., 2021; Remman et al., 2022; Schreiber et al., 2021; Theumer et al., 2022; Zhang et al., 2021), prior work has largely focused on specific applications, offering no formal analysis of what the resulting explanations actually mean. Without a clear theoretical foundation, these methods leave key questions unanswered: What does a Shapley value explain in reinforcement learning? How should it be interpreted? And how do previous explanations relate to one another? In contrast, SVERL provides a principled mathematical foundation for addressing these questions. It situates existing approaches within a unified perspective, revealing their underlying assumptions and limitations, and showing how they can yield incomplete, misleading, or easily misinterpreted explanations.
@@
The remainder of this paper is structured as follows. In Section 2, we provide useful background information on reinforcement learning and Shapley values. In Section 3, we introduce the motivation and foundational questions behind explaining reinforcement learning. In Sections 4, 5 and 6, we develop a unified framework for explaining agent behaviour, performance, and value estimation, respectively. In Section 7, we demonstrate the proposed framework in various illustrative domains. In Section 8, we discuss how the framework can be used, interpreted, and extended. We review related work in Section 9, propose directions for future research in Section 10, and conclude in Section 11.
@@ 2 Background
We indicate random variables by capital letters (e.g. $X$, $S$), their values by lowercase letters (e.g. $x$, $s$) and their domains with calligraphic fonts (e.g. ${\cal X}$, ${\cal S}$). Functions are also denoted by lowercase letters (e.g. $g$); it will be clear from context whether a lowercase letter refers to a function or a value. Random variables are assumed to be vectors, with ${\cal X}_{i}$ indicating the domain for value $x_{i}$ at index $i$. Given a base set =${\cal F}=\{1,2,\dots,n\}$, where $n$ is the size of the domain ${\cal X}$, we use $\bar{\mathcal{C}}\equiv{\cal F}\setminus\mathcal{C}$ to represent a set’s complement, and $x_{\mathcal{C}}\equiv\{x_{i}:i\in\mathcal{C}\}$ to denote a subset of values for $\mathcal{C}\subseteq{\cal F}$. Additionally, we use $\Delta({\cal A})$ to denote the probability simplex over the domain ${\cal A}$. That is, the function $\rho:S\rightarrow\Delta({\cal A})$ expresses a probability mass function $\rho(\cdot|s)$ over ${\cal A}$, for each $s\in{\cal S}$.
@@ 2.1 Reinforcement Learning
Reinforcement learning is an approach to artificial intelligence that models the interaction of an agent with its environment as the agent seeks to achieve desired outcomes. We follow the common approach of modelling this interaction as a Markov Decision Process (MDP), defined as a 5-tuple $({\cal S},{\cal A},p,r,\gamma)$. The environment is initiated in state $S_{0}\in{\cal S}$, according to some initial state distribution $d(s)\coloneqq\text{Pr}(S_{0}=s)$. At decision step $t$, $t\geq 0$, the agent observes the current state of the environment, $S_{t}\in{\cal S}$, and executes action $A_{t}\in{\cal A}(s_{t})$, where ${\cal A}(s_{t})$ is the set of executable actions in $s_{t}$. Consequently, the environment transitions to a new state, $S_{t+1}\in{\cal S}$, according to the transition kernel $p(s^{\prime}\,|\,s,a)\coloneqq\text{Pr}(S_{t+1}=s^{\prime}\,|\,S_{t}=s,A_{t}=a)$, and returns a reward $R_{t+1}\in\mathbb{R}$, whose expectation for a transition from state $s$ to state $s^{\prime}$ under action $a$ is denoted $r(s,a,s^{\prime})=\mathbb{E}[R_{t+1}\,|\,S_{t}=s,A_{t}=a,S_{t+1}=s^{\prime}]$. The return $G_{t}$ from decision step $t$ is defined as the discounted sum of all future rewards:
@@
where $\gamma\in[0,1]$ is the discount factor.
We assume that states can be decomposed into features, $S=(S_{1},S_{2},\dots,S_{n})$, where each $S_{i}\in{\cal S}_{i}$ represents the value of an individual feature, such as the velocity of an autonomous vehicle. We will use these state features as the basic elements of our explanations of reinforcement learning.
A policy $\pi:{\cal S}\rightarrow\Delta({\cal A})$ maps each state to a probability distribution over actions. The agent’s objective is to learn an optimal policy that maximises the expected return from every state. Reinforcement learning methods commonly compute the optimal policy by alternating between policy evaluation: estimating the expected return from state $s$ when following policy $\pi$, $v^{\pi}(s)\coloneqq\mathbb{E}_{\pi}{\left[G_{t}\,|\,S_{t}=s\right]}$, or the expected return of executing action $a$, $q^{\pi}(s,a)\coloneqq\mathbb{E}_{\pi}\left[G_{t}\,|\,S_{t}=s,A_{t}=a\right]$, and policy improvement: deriving a new policy $\pi^{\prime}$ by acting greedily with respect to $q^{\pi}(s,a)$:
$\pi^{\prime}(s)\coloneqq\underset{a\in{\cal A}}{\operatorname*{arg\,max}}\,q^{\pi}(s,a).$
@@ 2.2 Shapley Values in Game Theory
In the parliamentary voting example, the Shapley values for the three parties are: $\phi_{\text{A}}=\frac{1}{3}$, $\phi_{\text{B}}=\frac{1}{3}$, $\phi_{\text{C}}=\frac{1}{3}$. Using the Shapley values, we can conclude that all three parties make the same contribution towards each vote and hence have the same voting power.
Shapley values are the unique solution to four axioms, described below, that specify the conditions that must be met for a fair allocation of contributions among the players of a coalitional game. While there are other contribution schemes in the literature, Shapley values offer the only contribution scheme derived from these axioms of fairness, accounting for interdependencies between players (often observed in practice).
Let $\phi_{i}\left(v\right)$ denote how much player $i$ contributes to the outcome of a coalitional game $({\cal F},v)$. For fair allocation of credit, $\phi_{i}\left(v\right)$, $\forall i$, must satisfy:
  • Axiom 1 (Efficiency). The difference in value between the outcome of the game, $v({\cal F})$, and the fixed payoff for the game with no players, $v(\emptyset)$, is fully distributed over the set of players:
    $v({\cal F})-v(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(v)$
  • Axiom 2 (Linearity). If two independent coalitional games, $u$ and $v$, are combined into a new game with the characteristic function $\alpha u+\beta v$ for some scalars $\alpha,\beta\in\mathbb{R}$, then each player’s contribution in the combined game must be the weighted sum of their contributions in the individual games:
    $\phi_{i}(\alpha u+\beta v)=\alpha\phi_{i}(u)+\beta\phi_{i}(v)$
  • Axiom 3 (Symmetry). Two players are assigned the same contribution if they make equal marginal contributions to the outcome of the game for all possible coalitions of players:
    $\phi_{i}(v)=\phi_{j}(v)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ v(\mathcal{C}\cup\{i\})=v(\mathcal{C}\cup\{j\})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\,\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}$
  • Axiom 4 (Nullity). A player is assigned zero contribution if they make zero marginal contribution to the outcome of the game for all possible coalitions of players:
    $\phi_{i}(v)=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ v(\mathcal{C}\cup\{i\})=v(\mathcal{C})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\,\mathcal{C}\subseteq{\cal F}\setminus\{i\}$
The seminal paper by Shapley (1953) proved that Shapley values provide the unique solution that satisfies these four axioms. [^1] The efficiency, linearity, and symmetry axioms are sufficient to derive the Shapley value, while the nullity axiom is a direct consequence of them.
@@ 2.3 Explaining Supervised Learning with Shapley Values
At the specific input point $x$, the prediction $f(x)$ is made collectively by the feature values $x_{1},x_{2},\dots,x_{n}$. This prediction can be modelled as a coalitional game $\text{G}=({\cal F},f_{x})$, where the players ${\cal F}$ are the feature values $x$. The outcome of the game is the prediction $f(x)$. The characteristic function is $f_{x}:2^{|{\cal F}|}\rightarrow{\cal Y}$, where $f_{x}(\mathcal{C})$ returns the prediction at $x$ when the values of features not in $\mathcal{C}\subseteq{\cal F}=\{1,2,\dots,n\}$ are unknown. The game quantifies how different subsets of feature values $(x_{1},x_{2},\dots,x_{n})$ contribute to the prediction $f(x)$; its Shapley values $\phi_{1}(f_{x}),\dots,\phi_{n}(f_{x})$ uniquely satisfy four axioms defining a fair credit allocation:
  • Axiom 1 (Efficiency). The difference between the prediction when the values of all features are known, $f_{x}({\cal F})$, and the fixed prediction when the values of no features are known, $f_{x}(\emptyset)$, is fully distributed over the feature values:
    $f_{x}({\cal F})-f_{x}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(f_{x})$
  • Axiom 2 (Linearity). A feature value’s contribution to the prediction of a linear sum of independent models, $f$ and $g$, is equal to the sum of the feature value’s contributions towards the predictions of each model:
    $\phi_{i}(\alpha f_{x}+\beta g_{x})=\alpha\phi_{i}(f_{x})+\beta\phi_{i}(g_{x})$
  • Axiom 3 (Symmetry). Two feature values that make equal marginal contributions to the prediction of a model, when used (for prediction) with all possible subsets of feature values, are assigned the same contribution:
    $\phi_{i}(f_{x})=\phi_{j}(f_{x})\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ f_{x}(\mathcal{C}\cup\{i\})=f_{x}(\mathcal{C}\cup\{j\})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}$
  • Axiom 4 (Nullity). A feature value that makes zero marginal contribution to the prediction of a model, when used with all possible subsets of feature values, is assigned zero contribution:
    $\phi_{i}(f_{x})=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ f_{x}(\mathcal{C}\cup\{i\})=f_{x}(\mathcal{C})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}$
@@
Computing Shapley values requires making predictions $f_{x}(\mathcal{C})$ for all subsets of features $\mathcal{C}\subseteq{\cal F}$. However, the function $f(x):{\cal X}\rightarrow{\cal Y}$ specifies the prediction only when $\mathcal{C}={\cal F}$. It is therefore necessary to make some assumptions on how the model will predict when some features are absent. There are multiple approaches one can take. One common approach is to represent the unknown feature values by marginalising over their conditional distribution $p(X_{\bar{\mathcal{C}}}\,|\,X_{\mathcal{C}}=x_{\mathcal{C}})$ (Štrumbelj and Kononenko, 2014; Lundberg and Lee, 2017). Another common approach is to marginalise over their joint marginal distribution $p(X_{\bar{\mathcal{C}}})$ (Štrumbelj and Kononenko, 2010). We briefly explain these two approaches below.
@@
As a consequence of assuming features are independent, marginal characteristic functions consider impossible feature value combinations when features are dependent (Frye et al., 2020a). They were once treated as a relatively easy approximation of conditional characteristic functions. However, it has been argued that marginal characteristic functions may be a better way to represent unknown feature values because the resulting Shapley value for a feature not explicitly used by a model is always zero; if using the conditional characteristic function, the Shapley value for such a feature may be non-zero (Sundararajan and Najmi, 2020; Janzing et al., 2020).
For example, the function $f(X_{1},X_{2})=X_{1}$ never uses the feature $X_{2}$. However, if $X_{2}$ depends on $X_{1}$ (or is correlated with $X_{1}$), using the conditional characteristic function results in non-zero Shapley values for $x_{2}$ at any point $x$. These Shapley values “reflect the model’s dependence on the information communicated by each feature [value], rather than its algebraic dependence” (Jethani et al., 2021). When using the marginal characteristic function, on the other hand, the Shapley values do not account for any relationship between $x_{1}$ and $x_{2}$ because they assume that the features are independent. In other words, with the marginal approach, the Shapley value for $x_{2}$ is zero: $\phi(x_{2})=0$.

### v2

@@ Abstract
Reinforcement learning agents can achieve super-human performance in complex decision-making tasks, but their behaviour is often difficult to understand and explain. This lack of explanation limits deployment, especially in safety-critical settings where understanding and trust are essential. We identify three core explanatory targets that together provide a comprehensive view of reinforcement learning agents: behaviour, outcomes, and predictions. We develop a unified theoretical framework for explaining these three elements of reinforcement learning agents through the influence of individual features that the agent observes in its environment. We derive feature influences by using Shapley values, which collectively and uniquely satisfy a set of well-motivated axioms for fair and consistent credit assignment. The proposed approach, Shapley Values for Explaining Reinforcement Learning (SVERL), provides a single theoretical framework to comprehensively and meaningfully explain reinforcement learning agents. It yields explanations with precise semantics that are not only interpretable but also mathematically justified, enabling us to identify and correct conceptual issues in prior explanations. Through illustrative examples, we show how SVERL produces useful, intuitive explanations of agent behaviour, outcomes, and predictions, which are not apparent from observing agent behaviour alone.
Keywords: reinforcement learning, Shapley values, explainable artificial intelligence
@@ 1 Introduction
Reinforcement learning agents learn how to act through interaction with their environment, by taking actions, observing their consequences, and receiving scalar reward signals as feedback. They have achieved remarkable performance across diverse domains, including mastering games (Mnih et al., 2015; Silver et al., 2016, 2017; Vinyals et al., 2019; Schrittwieser et al., 2020; Wurman et al., 2022) and controlling complex physical systems (Bellemare et al., 2020; Degrave et al., 2022; Seo et al., 2024). Despite these successes, reinforcement learning agents remain difficult to interpret: they learn how to act but not how to explain themselves. This lack of explanation hinders deployment, particularly in safety-critical settings where accountability, oversight, and trust are essential. As one example, if a human operator cannot understand an agent’s decisions, should they entrust it with control over a nuclear fusion reactor (Degrave et al., 2022; Seo et al., 2024)?
In this paper, we identify three explanatory targets for reinforcement learning agents that together provide a comprehensive and informative view of the agent. These explanatory targets are behaviour, outcomes, and predictions. Behaviour refers to how an agent acts, that is, the decision to take a particular action in a given situation. Outcomes refer to the consequences of those actions, for example, the reward collected from the environment. Predictions refer to estimates of those outcomes, in other words, any predictions one may want to make about the future interaction of the agent with its environment. Each explanatory target captures a distinct and meaningful aspect of the agent-environment interaction. Together, they enable a comprehensive and structured understanding of the agent.
To explain these three explanatory targets—behaviour, outcomes, and predictions—we propose a unified framework that attributes responsibility to individual state features, which are numerical values that describe what the agent observes in its environment. We derive these attributions by using Shapley values (Shapley, 1953), which collectively satisfy a set of well-motivated axioms for fair and consistent credit assignment. This approach yields explanations that are not only interpretable but also mathematically justified and uniquely determined by the underlying axioms. We call the proposed framework Shapley Values for Explaining Reinforcement Learning (SVERL). Through illustrative examples, we show how SVERL produces intuitive explanations that improve our understanding of agent behaviour, outcomes, and predictions.
Although Shapley values have previously been applied to explain reinforcement learning agents (Rizzo et al., 2019; Carbone, 2020; He et al., 2021; Wang et al., 2020; Zhang et al., 2020; Liessner et al., 2021; Løver et al., 2021; Remman et al., 2022; Schreiber et al., 2021; Theumer et al., 2022; Zhang et al., 2021), existing work focuses on individual application domains and lacks a theoretical foundation. They often leave ambiguous what is being explained—conflating behaviour, outcomes, and predictions—or failing to distinguish them altogether. We show that without this clarity, explanations can be ambiguous, misleading, or incorrect. Critically, we show that earlier approaches do not explain outcomes, perhaps the most important and distinguishing aspect of reinforcement learning agents compared to explanations of supervised learning. SVERL explicitly distinguishes among explanatory targets and derives each explanation through its own axiomatic framework, ensuring that every attribution is precise, interpretable, and mathematically grounded. We show how existing approaches emerge as special cases of the general theoretical framework developed here, revealing how previous work relates to each other, what assumptions they make, and where their limitations lie.
In addition to Shapley-based methods, previous work has explained reinforcement learning agents by exploiting knowledge about their environment to analyse their learning dynamics (Amir et al., 2018; Huang et al., 2018; Rupprecht et al., 2019; Cruz et al., 2019; Juozapaitis et al., 2019; Madumal et al., 2020; Cruz et al., 2023); using simpler, more interpretable agent models to trade expressivity for transparency (Bastani et al., 2018; Jhunjhunwala, 2019; Silva et al., 2020; Topin et al., 2021; Bewley and Lawry, 2021); and attributing feature influence with alternative methods that lack the theoretical guarentees of Shapley values (Wang et al., 2016; Greydanus et al., 2018; Puri et al., 2020). In contrast, SVERL provides explanations that are theoretically grounded and generally applicable, without assumptions about models, dynamics, or environment structure.
While our focus is on establishing the theoretical foundations of SVERL, we also explore some of its natural extensions. Because it inherits the foundations of Shapley values, SVERL can naturally incorporate developments from the broader attribution literature. We illustrate this by showing, for example, how its explanations can be approximated when exact computation is impractical, enabling use in real-world settings, and how the framework might support explanations that extend beyond individual states to offer a broader view of agent behaviour, outcomes, and predictions.
These contributions position SVERL as a principled foundation for producing well-defined, interpretable, and comprehensive explanations of reinforcement learning agents.
@@
The remainder of this paper is structured as follows. In Section 2, we provide an overview of reinforcement learning and Shapley values. In Section 3, we introduce the motivation and foundational questions behind explaining reinforcement learning agents. In Sections 4, 5 and 6, we develop a unified framework for explaining agent behaviour, outcomes, and predictions. In Section 7, we present illustrative examples. In Section 8, we discuss how the framework can be interpreted and used in practice. We review related work in Section 9, propose directions for future research in Section 10, and conclude in Section 11.
@@ 2 Background
We indicate random variables by capital letters (e.g. $X$, $S$), their values by lowercase letters (e.g. $x$, $s$) and their domains with calligraphic fonts (e.g. ${\cal X}$, ${\cal S}$). Functions are also denoted by lowercase letters (e.g. $g$); it will be clear from context whether a lowercase letter refers to a function or a value. Random variables are assumed to be vectors, with ${\cal X}_{i}$ indicating the domain for value $x_{i}$ at index $i$. Given a base set =${\cal F}=\{1,2,\dots,n\}$, where $n$ is the dimension of the domain ${\cal X}$, we use $\bar{\mathcal{C}}\equiv{\cal F}\setminus\mathcal{C}$ to represent a set’s complement, and $x_{\mathcal{C}}\equiv\{x_{i}:i\in\mathcal{C}\}$ to denote a subset of values for $\mathcal{C}\subseteq{\cal F}$. Additionally, we use $\Delta({\cal A})$ to denote the probability simplex over the domain ${\cal A}$. That is, the function $\rho:S\rightarrow\Delta({\cal A})$ expresses a probability mass function $\rho(\cdot|s)$ over ${\cal A}$, for each $s\in{\cal S}$.
@@ 2.1 Reinforcement Learning
Reinforcement learning is an approach to artificial intelligence that models the interaction of an agent with its environment as the agent seeks to achieve desired outcomes. We follow the common approach of modelling this interaction as a Markov Decision Process (MDP), defined as a 5-tuple $({\cal S},{\cal A},p,r,\gamma)$. The environment is initiated in state $S_{0}\in{\cal S}$, according to some initial state distribution $d(s)\coloneqq\text{Pr}(S_{0}=s)$. At decision stage $t$, $t\geq 0$, the agent observes the current state of the environment, $S_{t}\in{\cal S}$, and executes action $A_{t}\in{\cal A}(s_{t})$, where ${\cal A}(s_{t})$ is the set of executable actions in $s_{t}$. Consequently, the environment transitions to a new state, $S_{t+1}\in{\cal S}$, according to the transition kernel $p(s^{\prime}\,|\,s,a)\coloneqq\text{Pr}(S_{t+1}=s^{\prime}\,|\,S_{t}=s,A_{t}=a)$, and returns a reward $R_{t+1}\in\mathbb{R}$, whose expectation for a transition from state $s$ to state $s^{\prime}$ under action $a$ is denoted $r(s,a,s^{\prime})=\mathbb{E}[R_{t+1}\,|\,S_{t}=s,A_{t}=a,S_{t+1}=s^{\prime}]$. The return $G_{t}$ from decision stage $t$ is defined as the discounted sum of all future rewards:
@@
where $\gamma\in[0,1]$ is a discount factor that weights the relative importance of future rewards.
We assume that states can be decomposed into features, $S=(S_{1},S_{2},\dots,S_{n})$, where each $S_{i}\in{\cal S}_{i}$ represents the (random) value of an individual feature, such as the velocity of an autonomous vehicle. We will use these state features as the basic elements of our explanations of reinforcement learning.
A policy $\pi:{\cal S}\rightarrow\Delta({\cal A})$ maps each state to a probability distribution over actions. The agent’s objective is to learn an optimal policy that maximises the expected return from every state. Reinforcement learning methods commonly compute the optimal policy by alternating between policy evaluation: estimating the expected return from state $s$ when following policy $\pi$, $v^{\pi}(s)\coloneqq\mathbb{E}_{\pi}{\left[G_{t}\,|\,S_{t}=s\right]}$, or the expected return of executing action $a$, $q^{\pi}(s,a)\coloneqq\mathbb{E}_{\pi}\left[G_{t}\,|\,S_{t}=s,A_{t}=a\right]$, and policy improvement: deriving a new policy by acting greedily with respect to $q^{\pi}(s,a)$:
$\underset{a\in{\cal A}}{\operatorname*{arg\,max}}\,q^{\pi}(s,a).$
@@ 2.2 Shapley Values in Game Theory
In the parliamentary voting example, the Shapley values for the three parties are: $\phi_{\text{A}}=\frac{1}{3}$, $\phi_{\text{B}}=\frac{1}{3}$, $\phi_{\text{C}}=\frac{1}{3}$. Using the Shapley values, we can conclude that, even though parties A and B both have 49 representatives, and party C has 2 representatives, all three parties make the same contribution towards each vote and hence have the same voting power.
Shapley values are the unique solution to four axioms, described below, that specify the conditions that must be met for a fair allocation of contributions among the players of a coalitional game. While there are other contribution schemes in the literature (such as Banzhaf values; Banzhaf III, 1964), Shapley values offer the only contribution scheme derived from these axioms of fairness.
Let $\phi_{i}\left(v\right)$ denote how much player $i$ contributes to the outcome of a coalitional game $({\cal F},v)$. For fair allocation of credit, the Shapley values $\phi_{i}\left(v\right)$, $\forall i$, must satisfy:
Axiom 1 (Efficiency)
The difference in value between the outcome of the game, $v({\cal F})$, and the fixed payoff for the game with no players, $v(\emptyset)$, is fully distributed over the set of players:
$v({\cal F})-v(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(v).$
Axiom 2 (Linearity)
If two independent coalitional games, $u$ and $v$, are combined into a new game with the characteristic function $\alpha u+\beta v$ for some scalars $\alpha,\beta\in\mathbb{R}$, then each player’s contribution in the combined game must be the weighted sum of their contributions in the individual games:
$\phi_{i}(\alpha u+\beta v)=\alpha\phi_{i}(u)+\beta\phi_{i}(v).$
Axiom 3 (Symmetry)
Two players are assigned the same contribution if they make equal marginal contributions to the outcome of the game for all possible coalitions of players:
$\phi_{i}(v)=\phi_{j}(v)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ v(\mathcal{C}\cup\{i\})=v(\mathcal{C}\cup\{j\})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\,\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}.$
Axiom 4 (Nullity)
A player is assigned zero contribution if they make zero marginal contribution to the outcome of the game for all possible coalitions of players:
$\phi_{i}(v)=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ v(\mathcal{C}\cup\{i\})=v(\mathcal{C})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\,\mathcal{C}\subseteq{\cal F}\setminus\{i\}.$
The seminal paper by Shapley (1953) proved that Shapley values are the unique solution that satisfies these four axioms. The efficiency, linearity, and symmetry axioms are sufficient to derive the Shapley value, while the nullity axiom is a direct consequence of them.
@@ 2.3 Explaining Supervised Learning with Shapley Values
At the specific input point $x$, the prediction $f(x)$ is made collectively by the feature values $x_{1},x_{2},\dots,x_{n}$. This prediction can be modelled as a coalitional game $({\cal F},f_{x})$, where the players ${\cal F}$ are the feature values $x$. The outcome of the game is the prediction $f(x)$. The characteristic function is $f_{x}:2^{|{\cal F}|}\rightarrow{\cal Y}$, where $f_{x}(\mathcal{C})$ returns the prediction at $x$ when the values of features not in $\mathcal{C}\subseteq{\cal F}=\{1,2,\dots,n\}$ are unknown. The game quantifies how different subsets of feature values $(x_{1},x_{2},\dots,x_{n})$ contribute to the prediction $f(x)$; its Shapley values $\phi_{1}(f_{x}),\dots,\phi_{n}(f_{x})$ uniquely satisfy four axioms defining a fair credit allocation, detailed below. When referring to a feature’s contribution, we implicitly mean the contribution of its value $x_{i}$ at input $x$.
Axiom 5 (Efficiency)
The difference between the prediction when the values of all features are known, $f_{x}({\cal F})$, and the fixed prediction when the values of no features are known, $f_{x}(\emptyset)$, is fully distributed over the features:
$f_{x}({\cal F})-f_{x}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(f_{x}).$
Axiom 6 (Linearity)
A feature’s contribution to the prediction of a linear sum of independent models, $f$ and $g$, is equal to the sum of the feature’s contributions towards the predictions of each model:
$\phi_{i}(\alpha f_{x}+\beta g_{x})=\alpha\phi_{i}(f_{x})+\beta\phi_{i}(g_{x}).$
Axiom 7 (Symmetry)
Two features that make equal marginal contributions to the prediction of a model, when used (for prediction) with all possible subsets of features, are assigned the same contribution:
$\phi_{i}(f_{x})=\phi_{j}(f_{x})\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ f_{x}(\mathcal{C}\cup\{i\})=f_{x}(\mathcal{C}\cup\{j\})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}.$
Axiom 8 (Nullity)
A feature that makes zero marginal contribution to the prediction of a model, when used with all possible subsets of features, is assigned zero contribution:
$\phi_{i}(f_{x})=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ f_{x}(\mathcal{C}\cup\{i\})=f_{x}(\mathcal{C})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}.$
@@
Computing Shapley values requires making predictions $f_{x}(\mathcal{C})$ for all subsets of features $\mathcal{C}\subseteq{\cal F}$. However, the function $f(x):{\cal X}\rightarrow{\cal Y}$ specifies the prediction only when $\mathcal{C}={\cal F}$. It is therefore necessary to make some assumptions on how the model will predict when some features are unknown. There are multiple approaches one can take. One common approach is to represent the unknown feature values by marginalising over their conditional distribution $p(X_{\bar{\mathcal{C}}}\,|\,X_{\mathcal{C}}=x_{\mathcal{C}})$ (Štrumbelj and Kononenko, 2014; Lundberg and Lee, 2017). Another common approach is to marginalise over their joint marginal distribution $p(X_{\bar{\mathcal{C}}})$ (Štrumbelj and Kononenko, 2010). We briefly explain these two approaches below.
@@
As a consequence of assuming features are independent, marginal characteristic functions consider impossible feature value combinations when features are dependent (Frye et al., 2020a). They were once treated as a relatively easy approximation of conditional characteristic functions. However, it has since been argued that marginal characteristic functions may be a better way to represent unknown feature values because the resulting Shapley value for a feature not explicitly used by a model is always zero; if using the conditional characteristic function, the Shapley value for such a feature may be non-zero (Sundararajan and Najmi, 2020; Janzing et al., 2020).
For example, the function $f(X_{1},X_{2})=X_{1}$ never uses the feature $X_{2}$. However, if $X_{2}$ depends on $X_{1}$ (or is correlated with $X_{1}$), using the conditional characteristic function results in non-zero Shapley values for $x_{2}$ at any point $x$. Shapley values using the conditional characteristic function “reflect the model’s dependence on the information communicated by each feature [value], rather than its algebraic dependence” (Jethani et al., 2021). When using the marginal characteristic function, on the other hand, the Shapley values do not account for any relationship between $x_{1}$ and $x_{2}$ because they assume that the features are independent. In other words, with the marginal approach, the Shapley value for $x_{2}$ is zero: $\phi(x_{2})=0$.